A capillary tube with an inside diameter of πŸ“πŸ’πŸŽ ππ’Ž can support a 𝟏𝟎𝟎 π’Žπ’Ž column of liquid that has a density of πŸπŸ‘πŸ“πŸŽπŸŽ π’Œπ’ˆπ’Žβˆ’πŸ‘. If 𝒉 = 𝟎. 𝟏 then is it possible to find the surface tension of the liquid? If not, then write down the reason.



1
Expert's answer
2021-11-16T19:20:36-0500
"\sigma=0.25\rho ghr^2\\
\sigma=0.25(13500)(9.8)(0.1)(0.54\cdot0.001)^2\\=0.00096\frac{N}{m}"